OpenStudy (anonymous):

You deposit $7,500 into a savings plan at the end of every six months. The interest rate is 4% compounded semiannually. Find the value of the annuity after 4 years. Do not round until the final answer. Then, round to the nearest cent. A $210,912.00 B $439,372.27 C $55,757.13 D $64,372.27

OpenStudy (nottim):

While I don't remember completely, isn't this about annuities?

OpenStudy (anonymous):

Yes it is

OpenStudy (nottim):

So really, just plug this into the annuity formula.

OpenStudy (anonymous):

How many days is semi-annually?

OpenStudy (nottim):

Semi-annually is half a year. And half a year is 6 mths.

OpenStudy (anonymous):

so is R = 7500, r = 0.04, m =6/12 = 1/2, and t = 4?

OpenStudy (anonymous):

and then plug those numbers into the formula? R = [1+r/m]^mt - 1] over r/m?

OpenStudy (anonymous):

is that right?

OpenStudy (nottim):

I dunno. You're the actual student.